Which number(s) cannot be used in the following rational expression?
3x-9/a^2+10a+21

Answer by rwm(914)


3x-9/a^2+10a+21
a^2+10a+21 can not be zero
so find the roots of
a^2+10a+21
factor
(a+7)*(a+3)=0
and the roots are a=-7 and a=-3
check
49-70+21=0
9-30+21=0
so a cannot be -7 nor -3